OpenStudy (anonymous):

Neil took his family of five out to dinner at HogTastic Pit BBQ. They all ordered the buffet and drink at $12.80 each. Neil left a 15% gratuity for the waitress on the after-tax bill. Based on a tax rate of 7.5%, Neil ended up paying a total of $____ for dinner.

OpenStudy (imstuck):

12.80 * 5 = 64. 64 + (.075)64=68.80. That covers the tax of .075 times 64. Now figure the tip on that after-tax bill at .15(68.80) + 68.80=79.12

OpenStudy (imstuck):

Hopefully the dinner was worth it for $80!!!!!